# Parameter-free prediction of irradiation defect structures in tungsten at room temperature using stochastic cluster dynamics Sicong He1, Brandon Schwendeman2, George Tynan2, Jaime Marian1,3

### v1

@@ Abstract
The foundations of irradiation damage theory were laid in the 1950s and 60s within the framework of chemical reaction kinetics. While helpful to analyze qualitative aspects of irradiation damage, the theory contained gaps that delayed its implementation and applicability as a predictive tool. The advent of computer simulations with atomistic resolution in the 80s and 90s revealed a series of mechanisms that have proved essential to understand key aspects of irradiation damage in crystalline solids. However, we still lack a comprehensive model that can connect atomic-level defect physics with experimental measurements of quantitative features of the irradiated microstructure. In this work we present a mesoscale model that draws from our improved understanding of irradiation damage processes collected over the last few decades, bridging knowledge gained from our most sophisticated atomistic simulations with defect kinetics taking place over time scales many orders of magnitude larger than atomic interaction times. Importantly, the model contains no adjustable parameters, and combines several essential pieces of irradiation damage physics, each playing an irreplaceable role in the context of the full model, but of limited utility if considered in isolation. Crucially, we carry out a set of experiments carefully designed to isolate the key irradiation damage variables and facilitate validation. Using tungsten as a model material, we find exceptionally good agreement between our numerical predictions and experimental measurements of defect densities and defect cluster sizes.
@@ 1 Introduction
The study of irradiation damage and its accumulation and effects on material properties is an exceedingly complex problem that has attracted considerable attention over the last several decades [1, 2, 3, 4, 5, 6, 7]. Irradiation brings the system to a far-from-equilibrium state generally characterized by defect concentrations many orders of magnitude higher than in thermal conditions. This triggers a number of transformations that lead to microstructural evolution, understood as a series of processes by which the system changes its internal structure to reduce its defect concentrations and lower its free energy.
@@
A popular choice in the study of irradiation damage in metals is tungsten (W), which has attracted substantial attention in recent times as a candidate plasma-facing material in fusion energy devices [8, 9, 10, 11, 12, 13]. As such, considerable effort has been put on developing material models that can capture the fine-scale features of irradiation damage in W. Electronic structure calculations based on density functional theory (DFT) are typically employed to calculate fundamental material properties and defect energetics [14, 15, 16, 17], while classical molecular dynamics (MD) simulations are used to obtain transport properties and simulate collisional displacement events [18, 19, 20, 21]. However, while atomistic simulations provide high physical accuracy, they do so at the expense of computational efficiency, eliminating them as a viable tool to reach irradiation doses of interest in experimental studies. Instead, simulations of high-dose irradiation in pure W are done using ‘object’ kinetic Monte Carlo (okMC), or mean-field rate theory (MFRT) methods [14, 22, 23, 10, 24, 25, 26, 27, 28, 29, 30]. These methods make use of computationally or experimentally obtained defect properties to surge above atomic-level time dynamics and extend the simulations’ time reach into the dpa scale. Unfortunately, lacking a clear reference to benchmark against, these studies are hampered by the need to incorporate the large number of irradiation mechanisms that have been proposed for W [10]. Recent experiments have been carried out to fill this gap [31, 32, 33], providing much needed information about the defect species present in the microstructure as a function of key irradiation variables.
@@
However, given the breadth and depth of irradiation effects in W that are captured in the models, it is of interest to carry out targeted experiments to isolate specific variables that may help us discriminate among the key mechanisms controlling the material response in each case. The main objective of this work is to perform a set of carefully designed irradiation experiments to help shed light on some of the most fundamental response mechanisms of irradiated W. We intentionally use tungsten specimens with minimum internal microstructure, i.e., well-annealed, very high-purity W single crystals with no dislocations or grain boundaries. With this choice, we eliminate the need to consider empirical bias factors for defect absorption by sinks, which are traditionally a source of confusion in many studies of irradiation damage modeling, particularly in metals [34, 35, 29]. Our simulations and experiments are conducted at 290 K, below irradiation recovery Stage III in W, where vacancy motion is negligible. All together, these conditions present a unique opportunity to validate specific aspects of the physical model of irradiation damage in W.
@@ 2 Physical model of irradiation damage
Mean-field rate theory provides the basis within which a system of partial differential equations (PDE) representing the coupled cluster dynamics can be solved to simulate the evolution of irradiation damage in materials [36, 37, 38]. Figure 1 shows the generalized mean-field cluster dynamics PDE based on classical nucleation theory. As the figure illustrates, the equation is formulated as a linear superposition of process change rates, truncated to second-order defect interactions to simplify the many-body problem111This is generally a very good approximation, as the interaction cross sections for three-body reactions and above are extremely small [39].. Numerical integration of this PDE system is feasible on modern computers, although to obtain practicable solutions, it is still necessary to apply certain numerical workarounds to accelerate the simulations [40, 41].
@@
Models such as this one are useful because they are formulated in a general and flexible manner intended to accommodate our latest knowledge in irradiation effects coming from a number of more accurate simulation methods. These may include (i) correlated defect distributions from displacement cascades, (ii) irradiation phenomena such as radiation enhanced diffusion, (iii) radiation induced precipitation and segregation, (iv) multi-species chemistry, etc. This information is generally encoded in coefficient matrices with the same dimensionality as that of the the PDE system. However, the above system of equations is designed to capture the entire defect cluster population from , where is the largest cluster size, which can in some cases reach millions of point defects. As such, the dimensionality of the cluster population can be extremely large, making the solution of these systems a tremendous numerical challenge. This also means that the physical coefficient matrices can be very large, sparse in some regions of the cluster dimensional space, while concentrated in others, with elements that have widely disparate values, and which can be time-evolving. This represents a mathematical challenge that even the most formidable deterministic solvers can struggle with.
@@
One can however solve the PDE system within the mean-field approximation by treating each term in the coefficient matrix as if it was an event rate. In this fashion, one naturally introduces the stochastic variability of explicit stochastic methods such as the kinetic Monte Carlo (kMC) approach but the mathematical form of MFRT methods is preserved. Event rates must be defined for a specific material volume, much like for kMC simulations, but matrix coefficients are obtained as for MFRT simulations. Likewise, spatial derivatives (diffusion terms) can be converted to rates in the simulation volume using the divergence theorem with standard discretizations as in the deterministic case. We call this method stochastic cluster dynamics (SCD) [42]. A brief review of SCD is provided in A, and for the remainder of this paper, the SCD method should be considered a stochastic version of the MFRT method, identical in its theoretical basis but more versatile and flexible in terms of the domain of application of the model and the numerical robustness of its solutions.
@@ 2.1 Experimental
The experiments have been designed to prioritize validation and facilitate comparison with the model predictions. Specifically, experimental and simulation conditions are harmonized by taking into consideration the following points:
@@
Material microstructure: the coefficients shown in the second term in the r.h.s. of Fig. 1, known as sink strengths, are difficult to calculate and subject to a large degree of empiricism [35]. Using well-annealed, high-purity single crystals avoids the presence of grain boundaries, dislocations, and precipitates, typically the most potent defect sinks found in crystalline materials.
@@
Irradiation temperature: the irradiations were conducted at 290 K, which is below the Stage III temperature below which there is no vacancy mobility [43, 44]. This leaves single self-interstitial atoms (SIA) and di-interstitials as the sole mobile species in the system [45, 46, 43, 47, 48], which simplifies the kinetics in the simulations.
@@
Irradiation with high-energy W self-ions ensures sufficient penetration to isolate the high-damage region of the specimen from its front and back surfaces, eliminating spurious local effects resulting from undue proximity of defects to the free surfaces. As well, implanting with W self-ions eliminates potential chemical interactions between the ions and the host lattice.
The experimental data were acquired using X-ray diffuse scattering (XRDS) [31], which can probe macroscopic volumes and thus isolate the irradiated samples from surface effects, as is often the case in TEM foils [49]. As well, XRDS provides a higher defect size sensitivity, resolving defects as small as 0.5 nm (tens of atoms, well below typial TEM resolution limits) and can distinguish between defect types, i.e., vacancies and self-interstitials and clusters thereof, adding an extra opportunity for comparison with defect evolution models. However, note that, due to resolution disparities, XRDS does not necessarily capture equal numbers of vacancies and self-interstitials in the clusters. This is again an opportunity for verification of the model predictions with respect to the experimental cluster concentrations.
@@ 2.2 Simulation conditions
The starting point for simulating irradiation damage processes in crystals is the PKA (primary knock-on atom) energy distribution function, which describes the energies with which recoils are produced during irradiation. For ions, the PKA distribution is extracted from SRIM simulations in full cascade mode [51]. Figure 2 shows the cumulative PKA energy distribution functions, , for the experiments performed by Schwendeman et al. [31] in pure W. The SRIM calculations indicate that the energy expended on lattice damage out of the total 10.8 MeV of ion energy is 2.65 MeV. Additionally, the average number of displacements, , per ion track was 18,373. The mean PKA energy can then be obtained from Fig. 2 as:
which reflects the depth dependence of recoil production in ion irradiation experiments through the variable , extending from the front surface at to the back surface at , as well as its dependence on energy . For the present case, we obtain a value of keV, although it is important to note, as we will demonstrate below, that is not a representative substitute of the full PKA energy spectrum. Damage accumulates at a rate of dpas-1 222We use the common unit of displacements per atom or ‘dpa’ to represent the damage fluence. In our experiments, 1 dpa amounts to a dose of approximately ions per cm2 [31].. These, and the rest of the simulation parameters are given in Table 1.
The simulations also include a one-hour annealing period at 290 K. This annealing is simulated by allowing the defects present in the system after irradiation to evolve according to their thermal relaxation laws (see terms with exponential terms in Fig. 1 and eq. (5)). Thermal annealing is generally controlled by diffusion of mobile species and dissociation of large immobile species by emission of monomers.
@@ 2.3 Key physics ingredients of the model
The model captures several crucial pieces of irradiation defect physics without which it would have not been possible to match experimental measurements. Some of these have only recently become available in the form of very recent works:
@@
Using cascade defect distributions obtained directly from MD simulations, without accounting for post-cascade correlated recombination.
@@
Next, we discuss each one of these times in detail as they have been implemented in the model.
@@ 2.3.1 Defect production in high-energy displacement cascades
The quantity in eq. (5) represents the insertion rate of defect species in the system. Defects are introduced in discrete bursts, each representing an independent displacement cascade, which are generated using correlations extracted from extensive databases obtained with molecular dynamics (MD) simulations of high-energy displacement cascades in W [52, 53, 54]. The information contained in these databases includes (i) the total number of Frenkel pairs, , produced by a PKA with a given energy, (ii) the fraction of defects forming part of clusters, and (iii) the cluster size distribution. Here we use the following expressions
These correlations are all dynamically sampled by the SCD code in every iteration and defects are introduced according to the PKA energy distribution provided in Fig. 2. Note that some of these works are very recent (and supersede previous studies [55, 56]) and thus only available to the nuclear materials community in the last few months.

### v3

@@ Abstract
The foundations of irradiation damage theory were laid in the 1950s and 60s within the framework of chemical reaction kinetics. While helpful to analyze qualitative aspects of irradiation damage, the theory contained gaps that delayed its implementation and applicability as a predictive tool. The advent of computer simulations with atomistic resolution in the 80s and 90s revealed a series of mechanisms that have proved essential to understand key aspects of irradiation damage in crystalline solids. However, we still lack a comprehensive model that can connect atomic-level defect physics with experimental measurements of quantitative features of the irradiated microstructure. In this work, we present a mesoscale model that draws from our improved understanding of irradiation damage processes collected over the last few decades, bridging knowledge gained from our most sophisticated atomistic simulations with defect kinetics taking place over time scales many orders of magnitude larger than atomic interaction times. Importantly, the model contains no adjustable parameters and combines several essential pieces of irradiation damage physics: each playing an irreplaceable role in the context of the full model but of limited utility if considered in isolation. Crucially, we carry out a set of experiments carefully designed to isolate the key irradiation damage variables and facilitate validation. Using tungsten as a model material, we find exceptionally good agreement between our numerical predictions and experimental measurements of defect densities and defect cluster sizes.
@@ 1 Introduction
The study of irradiation damage in crystalline materials and its accumulation and effects on material properties is an exceedingly complex problem that has attracted considerable attention over the last several decades weber2000models, woo2005modeling, wirth2007does, becquart2011modeling, dudarev2013density, xiao2020theoretical, fan2025atomistic. Irradiation brings the system to a far-from-equilibrium state generally characterized by defect concentrations many orders of magnitude higher than in thermal conditions. This triggers a number of transformations that lead to microstructural evolution, understood as a series of processes by which the system changes its internal structure to reduce its defect concentrations and lower its free energy.
@@
A popular choice in the study of irradiation damage in metals is tungsten (W), which has attracted substantial attention over the last decade as a candidate plasma-facing material in fusion energy devices rieth2013recent, rieth2025tungsten, marian2017recent, YU2019297, qian2021using, mcelfresh2024fracture. As such, considerable effort has been put on developing material models that can capture the fine-scale features of irradiation damage in W. Electronic structure calculations based on density functional theory (DFT) are typically employed to calculate fundamental material properties and defect energetics becquart2010microstructural, chen2012stability, boisse2014modelling, ventelon2012ab, while classical molecular dynamics (MD) simulations are used to obtain transport properties and simulate collisional displacement events fikar2009molecular, nandipati2015displacement, liu2023large, liu2025utilizing. Importantly, while atomistic simulations provide high physical accuracy, they do so at the expense of computational efficiency, eliminating them as a viable tool to reach irradiation doses of interest in experimental studies. Instead, simulations of high-dose irradiation in pure W are done using ‘object’ kinetic Monte Carlo (okMC), or mean-field rate theory (MFRT) methods becquart2010microstructural, MARIAN2012293, castin2017onset, marian2017recent, ZHANG2023101443, jin2018breaking, castin2018object, mason2019atomistic, hou2021influence, ma2024initial, wu2025influence, mohamed2025investigation. These methods make use of computationally or experimentally obtained defect properties to surge above atomic-level time dynamics and extend the simulations’ time reach near or above 1 dpa. Unfortunately, lacking clear experimental references to benchmark against, these studies are hampered by the need to incorporate the large number of irradiation mechanisms that have been proposed for W marian2017recent. Fortunately, recent experimental studies have provided much needed information regarding the defect species present in the microstructure as a function of key irradiation variables schwendeman2025integral, hu2025new, zavavsnik2025microstructural, Schwendeman_2026, thereby providing new opportunities to refine and improve mesoscale irradiation models.
@@
Given the breadth and depth of irradiation effects in W that are captured in the models, it is of interest to carry out targeted experiments to isolate specific variables that may help us discriminate among the key mechanisms controlling the material response in each case. Accordingly, the main objective of this work is to perform a set of carefully designed irradiation experiments to help shed light on some of the most fundamental response mechanisms of irradiated W. We intentionally use W specimens with minimal internal microstructure: specifically, well-annealed, very high-purity W single crystals with negligible intrinsic defect densities (such as dislocations) and low mosaicity (i.e., low-angle grain boundaries). This eliminates the need to consider empirical bias factors for defect absorption by sinks, which are traditionally a source of uncertainty in many studies of irradiation damage modeling, particularly in metals wolfer1975stress, malerba2007object, wu2025influence. Furthermore, this enables X-ray diffuse scattering (XRDS) measurements schwendeman2025integral, Schwendeman_2026 to be used to quantify the irradiation-induced defect cluster densities and size distributions in the material, including clusters with sizes below the typical resolution limits of transmission electron microscopy (TEM) measurements, for experimental validation. Our simulations and experiments are also conducted at 290 K, below irradiation recovery Stage III in W, where vacancy motion is negligible. Collectively, these conditions present a unique opportunity to validate aspects of the physical model of irradiation damage in W pertaining specifically to the cascade-induced defect cluster distributions and the subsequent room-temperature thermal evolution of this defect population.
@@ 2 Physical model of irradiation damage
Mean-field rate theory provides the basis within which a system of partial differential equations (PDEs) representing the coupled cluster dynamics can be solved to simulate the evolution of irradiation damage in materials golubov2007kinetics, ghoniem2020rate, odette1988mechanisms. Figure 1 shows the generalized mean-field cluster dynamics PDE based on classical nucleation theory. As the figure illustrates, the equation is formulated as a linear superposition of process change rates, truncated to second-order defect interactions to simplify the many-body problem111This is generally a very good approximation, as the interaction cross sections for three-body reactions and above are extremely small korobeinikov1975calculation.. Numerical integration of this PDE system is feasible on modern computers, although to obtain practicable solutions, it is still necessary to apply certain numerical workarounds to accelerate the simulations xu2020grouping, cao2006efficient.
@@
Models as described in Fig. 1 are useful because they are formulated in a general and flexible manner intended to accommodate our latest knowledge in irradiation effects coming from a number of more accurate simulation methods. These may include (i) correlated defect distributions from displacement cascades, (ii) irradiation phenomena such as radiation enhanced diffusion, (iii) radiation induced precipitation and segregation, (iv) multi-species chemistry, etc. This information is generally encoded in coefficient matrices with the same dimensionality as that of the the PDE system. However, the above system of equations is designed to capture the entire defect cluster population from , where is the largest cluster size, which can in some cases reach millions of point defects. As such, the dimensionality of the cluster population can be extremely large, leading to large, potentially time-evolving, physical coefficient matrices that are sparse in some regions of the cluster dimensional space while concentrated in others, with elements that have widely disparate values. This complexity presents a significant numerical challenge that even the most robust deterministic solvers often struggle to overcome.
@@
The PDE system within the mean-field approximation can be solved by treating each term in the coefficient matrix as an event rate. In this fashion, one naturally introduces the stochastic variability of explicit stochastic methods such as the kinetic Monte Carlo (kMC) approach but, crucially, the mathematical form of MFRT methods is preserved. Event rates must be defined for a specific material volume, similar to kMC simulations, but matrix coefficients are obtained just as in MFRT simulations. Likewise, spatial derivatives (diffusion terms) can be converted to rates in the simulation volume using the divergence theorem with standard discretizations as in the deterministic case. We call this method stochastic cluster dynamics (SCD) marian2011stochastic, and a brief review of SCD is provided in A. For the remainder of this paper, the SCD method should be regarded as a stochastic version of the MFRT method: identical in its theoretical basis but more versatile and flexible in terms of the domain of application of the model and the numerical robustness of its solutions.
@@ 2.1 Experimental
The experiments have been designed to facilitate comparison with the modeling predictions by considering the following points:
@@
Utilizing well-annealed, 99.999% purity W single crystals precludes the presence of high-angle grain boundaries and minimizes the densities of both low-angle grain boundaries and intrinsic defects (e.g., dislocations and precipitates). These are typically the most potent defect sinks found in crystalline materials, and thus their absence simplifies the treatment of defect interactions. In addition, this eliminates the need to compute the coefficients shown in the third term in the r.h.s. of Fig. 1, known as sink strengths, which are often difficult to calculate and subject to a large degree of empiricism malerba2007object.
@@
The irradiations were conducted at 290 K, which is below the Stage III temperature where vacancy mobility is effectively negligible. anand1978recovery, nambissan1992positron. This leaves single self-interstitial atoms (SIA) and di-interstitials as the sole mobile species in the system dicarlo1969stage, KEYS1970260, anand1978recovery, seidman1979point, wilson1980situ, which largely simplifies the defect kinetics in the simulations.
@@
Irradiation with 10.8 MeV W self-ions ensures sufficient penetration to isolate the high-damage region of the specimen from its front and back surfaces, thereby eliminating spurious local effects resulting from undue proximity of defects to the free surfaces. Moreover, implanting with W self-ions eliminates potential chemical interactions between the ions and the host lattice.
The irradiation-induced defect cluster densities and size distributions were experimentally quantified using ‘integral’ XRDS measurements schwendeman2025integral, Schwendeman_2026 around the surface-normal 110 Bragg reflection of the W single crystals. The analysis of these measurements is based on detailed modeling of the atomic structure and the lattice strain associated with irradiation-induced defects, as well as the diffuse scattering produced by statistical distributions of these defect clusters across macroscopic volumes of single crystals Dederichs_1971, Larson_1980, Ehrhart_1982, schwendeman2025integral, Schwendeman_dissertation, Schwendeman_2026. Importantly, XRDS is distinguished by its ability to determine the densities of defect clusters with sizes below the typical resolution limits of TEM measurements Larson_1987, Olsen_2016. In addition, due to the opposite signs of the long-range strain fields surrounding interstitial and vacancy-type defect clusters, XRDS can effectively extract separate size distributions for the two cluster types Larson_1987, Olsen_2016, Schwendeman_dissertation, Schwendeman_2026. These two capabilities, in particular, provide a level of detail regarding the defect size distributions produced directly by the displacement cascades as well as the distinct defect clustering fractions, binding energies, and thermal mobilities of interstitial and vacancy clusters that enable direct comparison with SCD modeling predictions.
XRDS measurements around Bragg reflections are primarily sensitive to defect clusters that produce strong strain fields in the surrounding lattice Ehrhart_1994, Larson_2019, Schwendeman_dissertation. In ion-irradiated W, TEM measurements YI2016105, Ferroni_2015, Guo_2020 and atomistic modeling byggmastar2025four have demonstrated that these will be interstitial and vacancy-type dislocation loops with primarily Burgers vectors. The XRDS measurements have been analyzed accordingly Schwendeman_2026 to determine separate size distributions for these two loop types with loop diameters ranging from to 40 nm, which are comprised of 7 to over 10,000 point defects. Importantly, existing experimental studies Hollingsworth_2022, zavavsnik2025microstructural and modeling Granberg_2021 have demonstrated that isolated vacancies (i.e. monovacancies) and very small globular (i.e. more spherically shaped) vacancy clusters are also produced under the irradiation conditions studied here (this is consistent with the experimental and modeling results presented in this work). These monovacancies and small vacancy clusters result in comparatively weak strain fields, and thus their densities and size distributions are not determined by the XRDS measurements.
@@ 2.2 Simulation conditions
The starting point for simulating irradiation damage processes in solids is the PKA (primary knock-on atom) energy distribution function that describes the energies with which recoils are produced during particle collisions. For 10.8 MeV self-ion irradiations of pure W used in the experiments performed by Schwendeman and co-workers schwendeman2025integral, Schwendeman_2026, the PKA distribution has been extracted from SRIM simulations in full cascade mode SRIM. Figure 2 shows the cumulative PKA energy distribution function, , as a function of depth beneath the sample’s surface. The SRIM calculations indicate that the average energy expended on lattice damage out of the total 10.8 MeV of ion energy is 2.65 MeV. Additionally, the average number of displacements per ion track is =18,373. The mean PKA energy is then obtained from Fig. 2 as:
which reflects the depth dependence of recoil production in ion irradiation experiments through the variable , extending from the front surface at to the back surface at , as well as its dependence on recoil energy . For the present case, we obtain a value of keV, although it is important to note that is not a representative substitute of the full PKA energy spectrum, as we will demonstrate below. Damage accumulates at a rate of dpas-1 222We use the common unit of displacements per atom or ‘dpa’ to represent the damage fluence. In this study with 10.8 MeV self-ion irradiation of W, 1 dpa amounts to a dose of approximately ions per m2 schwendeman2025integral., and the rest of the simulation parameters are given in Table 1.
In the configuration used here, XRDS measurements are not depth-sensitive, and thus the XRDS results provide a depth-integrated picture of the defect size distributions. The relative defect density is assumed to follow the SRIM-calculated damage-depth profile illustrated in Fig. 2, and the defect densities reported by the experimental analysis correspond to the peak damage depth at nm schwendeman2025integral, Schwendeman_2026. Therefore, for cross-consistency between experiments and modeling, here we sample PKA energies from the entire function, i.e., first we sample the depth variable, , with a probability given by the dependence of , followed by an energy sampling of . In order to match the experimental conditions, the simulations also include a one-hour annealing period at 290 K. This annealing is simulated by allowing the defects present in the system after irradiation to evolve according to their thermal relaxation laws (see terms with Boltzmann factors in Fig. 1 and eq. (5)). Thermal annealing is generally controlled by diffusion of mobile species and dissociation of large immobile species by emission of monomers as described in the following subsection and the Appendix.
@@ 2.3 Key physics ingredients of the model
The model captures several crucial pieces of irradiation defect physics that have proved essential to reproduce experimental measurements. Some of these are based on newly-acquired understanding from very recent works:
@@
Using cascade-induced defect distributions obtained directly from MD simulations, which do not account for post-cascade correlated recombination.
@@
The following subsections discuss each of these items in detail as they have been implemented in the model.
@@ 2.3.1 Defect production in high-energy displacement cascades
The quantity in eq. (5) represents the insertion rate of defect species in the system. Defects are introduced in discrete bursts, each representing an independent displacement cascade, which are generated using correlations extracted from extensive databases produced by molecular dynamics (MD) simulations of high-energy displacement cascades in W setyawan2015displacement, ZHANG2023101443, marian2025computational, byggmastar2025four. The relevant information provided by these databases includes (i) the total number of Frenkel pairs, , produced by a PKA with a given energy, (ii) the fraction of defects that form part of clusters and (iii) the distribution of the size of the clusters. Here we use the following expressions:
These correlations are all dynamically sampled by the SCD code in every iteration and defects are introduced according to the PKA energy distribution provided in Fig. 2. Note that some of the aforementioned MD studies are very recent (and supersede previous studies troev2011simulation, sand2013high) and thus only available to the nuclear materials community in the last few months.